OpenStudy (anonymous):

((3a^2)(b^-1)/c)^-3 solve. All I come up with is wrong. ): Please help and explain.

OpenStudy (anonymous):

since (b^-1/c) has a negative exponent you bring it to the denominator. right?

OpenStudy (anonymous):

the exponent of (b^-1/c) is -3 <--negative

OpenStudy (anonymous):

I'd assume you meant this: \[((3a^2)(\frac{b ^{-1}}{c}))^{-3}=((3a^2)(\frac{1}{bc}))^{-3}=((\frac{3a^2}{bc}))^{-3}=\frac{1}{(\frac{3a^2}{bc})^3}=\frac{b^3c ^{3}}{27a ^{6}}\]